OpenStudy (anonymous):

Please help me? The lead level in a child’s body is considered to be dangerously high if it exceeds 30 micrograms. Children come into contact with lead from a variety of sources,but are particularly susceptible to exposure from eating paint from toys.A random sample of 1000 children living in public housing projects in a particular city revealed that 200 of them had dangerously levels of lead in their bodies.Estimate the proportion of children with dangerously high lead levels in this population.If there 20,000 children living,estimate the number of children with dangerously high lead levels

OpenStudy (anonymous):

Do you want a point estimate or interval estimate? A good point estimate is just 200/1000. For an interval estimate, you need to know what level of confidence you are to use in constructing the interval.

OpenStudy (anonymous):

point estimate

OpenStudy (anonymous):

.2 = 200/1000

OpenStudy (anonymous):

And if it is interval estimate? what will be the results? Thank's for your help.

OpenStudy (anonymous):

Generally, you need to know what level of confidence you want to use in constructing the interval. Since you're working with proportions, you would use the normal distribution. The interval would be contructed in this way:\[p \pm z _{\frac{ \alpha }{ 2 }}\times \sqrt{\frac{ p \left( 1-p \right) }{ n }}\] where p is the point estimate, n is the sample size and \[z_{\frac{ \alpha }{ 2 }}\] is obtained from th enormal distribution. This assumes an infinite population. If it's a finite population you would use the finite population correction factor.

OpenStudy (anonymous):

i cant understand you? would you please explain to me the answer in steps?